# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v11

@@ 5 Mathematical Library Functions
At each point $x$ with an input uncertainty $\delta x$, the result uncertainty is calculated by variance arithmetic. The corresponding error deviation is determined by sampling as follows:
@@
    For each $\tilde{x}$, use the corresponding library function to compute the value error as the difference between using $\tilde{x}$ and using $x$ as the input.
@@ 5.1 Exponential
Figure 13 demonstrates that the calculated uncertainties using Formula (2.26) align closely with the measured value deviations for $e^{x+\delta x}$. Consequently, all error deviations remain very close to $1$, even though both the uncertainty and the value deviations increase exponentially with $x$ and $\delta x$.
@@ 5.2 Logarithm
Because $\log(x)$ has a distributional zero at $x=0$ all $log(x\pm\delta x)$ values are rejected when $P(x)>1/5$. Figure 14 demonstrates that the uncertainties calculated using Formula (2.29) align closely with the measured value deviations for $\log(x+\delta x)$, and the resulting error deviations remain very close to $1$ except when $P(x)>0.20086$.
@@ 5.3 Power
Figure 15 demonstrates that the uncertainties calculated for $(1\pm\delta x)^{c}$ using Formula (2.35) fit closely with the measured value deviations for $(1+\delta x)^{c}$, with the error deviations remaining near 11. Figure 16 reveals that the error deviations of $\sin(x+\delta x)$ align close to 11 for all the input exponent $c$ and the input uncertainty $\delta x>10^{15}$.
@@ 5.4 Sine
Figure 17 demonstrate that the uncertainties calculated using Formula (2.32) correspond closely to the measured value deviations for $\sin(x+\delta x)$. It also reveals that $\delta^{2}\sin(x)$ has the same periodicity as $\sin(x)$:
@@ 5.5 Numerical Errors for Library Functions
The combined numerical error of the library functions $e^{x}$ and $\log(x)$ is evaluated as either $\log(e^{x})-x$ or $e^{\log(x)}-x$. Using either variance arithmetic or conventional floating-point library functions produce identical value errors. Figure 19 demonstrates the corresponding result uncertainties, which reach a minimum when the input is a 2’s fraction such as $1$ or $1/2$. Figure 19 also shows that $e^{\log(x)}$ has much less error than $\log(e^{x})$. For $\log(e^{x})-x$, the error deviation is $0.409$ when $|x|\leq 1$ or approaches zero otherwise. The reason for surprising small value errors for $e^{\log(x)}-x$ is not clear at the moment.
@@ 5.6 Summary
Formula (2.26), (2.29), (2.35), and (2.32) provide effective estimates for their respective uncertainties for library functions. With added noise larger than $10^{-15}$, ideal coverage is achieved unless near a distributional pole where the error deviation approaches 0. In other cases, proper coverage is attainable.
@@ 6.2 Adjugate Matrix
The difference between $\widetilde{\mathbf{M}}^{A}$ and $\mathbf{M}^{A}$ defines the Adjugate Test. Figure 21 shows that the uncertainty means increase exponentially with both the input noises and the matrix size; however, such linear increase is segmented into two areas at input precision $10^{-10}$. Figure 22 shows that the ideal coverage is achieved for the input precision except at matrix size $8$ and input precision $10^{-17}$. The existence of ideal coverage validates Formula (6.9).
@@ 6.3 Floating Point Rounding Errors
The significand of the conventional floating-point representation [9] has $53$-bit resolution. Figure 25 shows that the histogram of the normalized errors is Delta distributed for the matrix size less than $8$, because the adjugate matrix calculation involves about $8\times 6=48$ significand bits for a matrix size $7$. When the matrix size is $8$, $8\times 7=56$ significand bits are needed so rounding occurs, which results in non-Delta distribution in Figure 25. The rounding error is also the reason why only at matrix size 8 and input noise precision $10^{-10}$, error deviation is no longer $1$ in Figure 22.
@@ 6.5 Matrix Inversion
Figure 28: The linear correlation between the precision of a matrix determinant (y-axis) to its condition number (x-axis). The legend shows the size of the matrix, as well as the type of the matrix as Random for randomly generated matrix, and Hilbert as the Hilbert matrix.
@@ 7.1 Moving-Window Linear Regression Algorithm
In many applications, data set $Y_{j}$} represents an input data stream where $j$ represents the time index or sequence index. $Y_{j}$} is referred to as a time-series input, where $j$ corresponds to $X_{j}$. A moving window algorithm [13] is applied within a small window centered on each $j$. For each calculation window, $X_{j}=-H,-H+1\dots H-1,H$} where $H$ is an integer constant specifying window’s half width. This choice ensures $\sum_{j}X_{j}=0$, which simplifies Formula (7.1) and (7.2) into Formula (7.3) and (7.4), respectively [1]:
@@
The values of $(\alpha_{j},\beta_{j})$ can be derived from the previous values $(\alpha_{j-1},\beta_{j-1})$, allowing Formula (7.3) and (7.4) to be reformulated into the progressive moving-window calculation given by Formula (7.5) and (7.6), respectively [1]
@@ 7.2 Variance Adjustment
When the time series contains uncertainty, directly applying Formula (7.5) and (7.6) will lead to a loss of precision, as both formulas reuse each input multiple times, accumulating the variance of that input at every usage. To avoid this, $\alpha_{j}$ and $\beta_{j}$ should still be calculated progressively using Formula (7.6) and (7.5), respectively, while the variance should instead be computed using Formula (7.7) and (7.8), respectively. Notably, Formula (7.8) is no longer progressive.
$\delta^{2}\alpha_{j}$ $=\sum_{X=-H+1}^{H}(\delta Y_{j-H+X})^{2}=\delta^{2}\alpha_{j-1}+(\delta Y_{j})^{2}-(\delta Y_{j-2H})^{2};$ (7.7)
$\delta^{2}\beta_{j}$ $=\sum_{X=-H+1}^{H}X^{2}(\delta Y_{j-H+X})^{2};$ (7.8)
@@
The specified input uncertainty is consistently $0.2$. Normal noises with a deviation of $0.2$ are added to the slopes, except for the segment $j=10\dots 19$ where Normal noises with a deviation of $2$ are introduced, representing a $10$-fold actual value over the specified uncertainty.
Figure 29 also presents the result of moving window fitting of $\alpha+\beta\;Y$ versus the time index $j$. The fitted values of $\alpha$ and $\beta$ follow the expected behave with a characteristic delay of $H$ in $j$. When (7.3) and (7.4) are applied to compute the uncertainties for $\alpha$ and $\beta$, they both increase exponentially with the time index $j$. In contrast, when Formula (7.3) and (7.4) are used exclusively for value calculation, while Formula (7.7) and (7.8) are applied for variance computation, the resulting uncertainties of $\alpha$ and $\beta$ are $\frac{\delta Y}{\sqrt{2H+1}}$, and $\frac{\delta Y}{\sqrt{\frac{H(H+1)(2H+1)}{3}}}$. Both are less than the input uncertainty $\delta Y$, due to the averaging effect of the moving window.
@@ 7.3 Unspecified Input Error
To determine the error deviation of $\alpha$ and $\beta$, the fitting procedure is applied to multiple time-series data sets, each generated with independent noise realizations. Figure 30 illustrates the resulting error deviation as a function of the time index $j$, which remains close to $1$ except in the range $j=10\dots 19$ where the actual noise is $10$-times greater than the specified value. This observation indicates that a larger than $1$ error deviation may signal the presence of unspecified additional input errors beyond rounding errors, such as the numerical errors in mathematical library functions.
@@ 8.1 DFT (Discrete Fourier Transformation)
For each signal sequence $h[k]$, where $k=0,1\dots N-1$, and $N$ is a natural number, the DFT (discrete Fourier transformation) $H[n]$, for $n=0,1\dots N-1$, along with its inverse transformation, is given by Formula (8.1) and (8.2) [13], respectively. In these expressions, $k$ denotes the time index while $n$ represents the frequency index. The frequency index and time index are not necessarily related to time unit and frequency unit, respectively. The naming is just a convenient way to distinguish the two opposite domains in DFT: the waveform domain $h[k]$ versus the frequency domain $H[n]$.
@@ 8.2 FFT (Fast Fourier Transformation)
  • For each output, each input is used only once, so there is no dependency problem when decomposing FFT into arithmetic operations involving Formula (2.11), (2.12), (2.13), and (2.14).
  • When $L$ is large, the large volume of input and output data enables high-quality statistical analysis.
  • The computational complexity is proportional to $L$, because increasing $L$ by 1 adds one more step involving a sum of multiplications.
  • Each step in the forward transformation increases the variance by a factor of $2$, so the uncertainty mean increases with the FFT order $L$ as $\sqrt{2}^{L}$. Because the reverse transformation divides the result by $2^{L}$, its uncertainty mean decreases with $L$ as $\sqrt{1/2}^{L}$. The uncertainty mean for the roundtrip transformation is therefore $\sqrt{2}^{L}\times\sqrt{1/2}^{L}=1$.
  • Forward and reverse transformations are identical except for a sign difference, so they are essentially the same algorithm, and any observed difference arises purely from the input data.
@@
  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum where most values are 0, causing its uncertainties to grow faster.
  • The reverse transformation spreads the precise frequency-domain spectrum (with most values being 0) back into a time-domain sine or cosine signals, causing its uncertainties to grow more slowly.
8.3 Modeling Errors of DFT and FFT
Although mathematically self-consistent, by implying a periodic boundary condition in the time domain [1], DFT and its efficient implementation FFT are only approximate for the mathematically defined continuous Fourier transformation [1]. They show no modeling error only when the input signal frequency $f$ satisfies $f=j\frac{2\pi}{N}$, and exhibit different degrees of modeling errors otherwise, with the modeling errors peaking at $f=(j+\frac{1}{2})\frac{2\pi}{N}$. Because of these modeling errors, using DFT and FFT as the digital implementation of continuous Fourier transformation is questionable, even though such usage is ubiquitous, and fundamental to many areas of applied mathematics [13].
To avoid the modeling errors of FFT, only Formula (8.1) and (8.2) are used in this paper.
@@ 8.4 Testing Signals
  • Results from Sin and Cos signals are statistically indistinguishable.
  • Results from Sin signals at different frequencies are also statistically indistinguishable.
Therefore, the results for Sin and Cos signals at all frequencies are pooled together for statistical analysis, under the combined category Sin/Cos signals.
@@ 8.5 Trigonometric Library Errors
Formula (8.1) and (8.2) restrict the use of $\sin(x)$ and $\cos(x)$ to $x=2\pi j/2^{L}$, in which $L$ is the FFT order. To minimize numerical errors in computing $\sin(x)$, indexed sine can replace standard library sine functions:
@@
Such constructed indexed $\sin(x)$ is named as the Quart indexed sine function. In contrast, the direct application of library $\sin(x)$ is named as the Library sine function.
Because the Quart sine function obeys the symmetry and periodicity of sine function strictly, it is superior to the library function.
  • Figure 31 shows that the value difference between Quart and Library $\sin(x)$ and the Quart $\sin(x)$ increases roughly linearly with $|x|$.
  • Figure 32 shows the value difference between Quart and Library $\cos(x)/\sin(x)$ also increases roughly linearly with $|x|$, but they are $10^{2}$ times larger than those of $\sin(x)$. Therefore, the linear spectrum in Formula (8.3) may contain huge numerical errors using library sine functions.
For both sine functions, the uncertainty of each $\sin(x)$ is assumed to be its ULP. Because Quart sine function has minimal Taylor expansion error due to its minimal range of $x$ in library $\sin(x)$, its true numerical errors is probably less than ULP, therefore its uncertainty should be sightly overestimated. In contrast, Figure 31 shows that library sine function has underestimated uncertainties.
@@ 8.6 Using Quart Sine for Sin/Cos Signals
Figure 33: FFT spectrum of $\sin(2\pi 3j/2^{6})$ using Quart sine function after the forward transformation computed by variance arithmetic. The legend distinguishes between the uncertainty and the value errors. The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
Figure 34: FFT waveform of $\sin(2\pi 3j/2^{6})$ using the Quart sine function after the reverse transformation computed by variance arithmetic. The legend distinguishes between the uncertainty and the value errors. The x-axis represents the time index, while the y-axis represents both the uncertainty and the value error.
Figure 35: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal at frequency 33 as a function of FFT order (x axis) and transformations (legend) using Quart sine function.
Figure 36: Histograms of normalized errors of Sin/Cos signals for forward, reverse and roundtrip transformations (legend) using Quart sine function. The FFT order is $18$.
Using Quart sine function, for a sine wave with $3$ as frequency, Figure 33 displays the spectrum for the forward transformation, while Figure 34 presents the waveform for the reverse transformation.
  • In the forward transformation, the value errors are slightly smaller than the corresponding result uncertainties, with an error deviation is $0.31$. In the reverse transformation, the error deviation is $0.43$. Both error deviations are less than $1$, confirming a slight overestimation of the sine uncertainty by ULP.
  • The uncertainty mean of the forward transformation is $7$ times larger than that of the reverse transformation.
@@
    When FFT order increases, the uncertainty mean of the forward transformation grows exponentially faster than that in the reverse transformation. At FFT order $18$, the uncertainty mean of the forward transformation is $10^{3}$ times more than that of the reverse transformation.
@@
    At FFT order $18$, Figure 36 shows that the normalized error distribution for the reverse transformation is wider than that for the forward transformation, but the two distributions do not differ by order-of-magnitude. Therefore, the value error grows at least $10^{2}$ times faster in the forward transformation than in the reverse transformation.
@@
    In forward transformation, error deviations reach their stable values quickly when FFT order $L\geq 4$. In contrast, in the reverse transformation, error deviations stabilize slowly with increasing FFT order until $L\geq 15$. The slow growth of normalized errors in the reverse transformation is attributed to its input data which contains all precise zeros except at the two frequency indexes.
Despite these differences, error deviations are between $0.2$ and $0.5$, suggesting that in variance arithmetic, uncertainties track value errors effectively in all cases.

### v12

@@ 6 Mathematical Library Functions
At each point $x$ with an input uncertainty $\delta x$, the result uncertainty is calculated by variance arithmetic. The corresponding error deviation is determined by sampling as:
@@
    For each $\tilde{x}$, use the corresponding library function to compute the value error as the difference between the outputs obtained using $\tilde{x}$ and $x$ as inputs.
@@ 6.1 Exponential
Figure 15 shows that the calculated uncertainties obtained using Formula (2.28) align closely with the measured value deviations for $e^{x+\delta x}$. Consequently, all error deviations remain very close to $1$, even though both the uncertainty and the value deviations increase exponentially with $x$ and $\delta x$.
@@ 6.2 Logarithm
Because $\log(x)$ has a distributional zero at $x=0$ all $log(x\pm\delta x)$ values are rejected when $P(x)>1/5$. Figure 16 shows that the uncertainties calculated using Formula (2.31) align closely with the measured value deviations for $\log(x+\delta x)$, and the resulting error deviations remain very close to $1$ except when $P(x)>0.20086$.
@@ 6.3 Power
Figure 17 shows that the uncertainties calculated for $(1\pm\delta x)^{c}$ using Formula (2.37) fit closely with the measured value deviations for $(1+\delta x)^{c}$, with the error deviations remaining near 11. Figure 18 reveals that the error deviations of $\sin(x+\delta x)$ remain close to 11 for all input exponents $c$ and input uncertainties $\delta x>10^{15}$.
@@ 6.4 Sine
Figure 19 shows that the uncertainties calculated using Formula (2.34) correspond closely to the measured value deviations for $\sin(x+\delta x)$. It also reveals that $\delta^{2}\sin(x)$ exhibits the same periodicity as $\sin(x)$:
@@ 6.5 Numerical Errors for Library Functions
The combined numerical error of the library functions $e^{x}$ and $\log(x)$ is evaluated as either $\log(e^{x})-x$ or $e^{\log(x)}-x$. Using either variance arithmetic or conventional floating-point library functions produces identical value errors. Figure 21 shows the corresponding result uncertainties, which reach a minimum when the input is a 2’s fraction such as $1$ or $1/2$. Figure 21 also shows that $e^{\log(x)}$ exhibits much smaller error than $\log(e^{x})$. For $\log(e^{x})-x$, the error deviation is $0.409$ when $|x|\leq 1$ or approaches zero otherwise. The reason for the unexpectedly small value errors for $e^{\log(x)}-x$ is not yet clear.
@@ 6.6 Summary
Formula (2.28), (2.31), (2.37), and (2.34) provide effective estimates of the corresponding library functions. When added noise exceeds $10^{-15}$ precision, ideal coverage is achieved except near a distributional pole where the error deviation approaches zero. In all other cases, proper coverage is attainable.
@@ 7.2 Adjugate Matrix
The difference between $\widetilde{\mathbf{M}}^{A}$ and $\mathbf{M}^{A}$ defines the Adjugate Test. As seen in Figure 23, the uncertainty means increase exponentially with both the input noises and the matrix size; however, such linear increase is segmented into two areas at input precision $10^{-10}$. Figure 24 shows that the ideal coverage is achieved for the input precision except at matrix size $8$ and input precision $10^{-17}$. The existence of ideal coverage validates Formula (7.9).
@@ 7.3 Floating Point Rounding Errors
The significance of the conventional floating-point representation [9] has a $53$-bit resolution. As shown in Figure 27, the histogram of the normalized errors is Delta distributed for the matrix size less than $8$, because the adjugate matrix calculation involves about $8\times 6=48$ significand bits for a matrix size $7$. When the matrix size is $8$, $8\times 7=56$ significand bits are needed so rounding occurs, which results in non-Delta distribution in Figure 27. The rounding error is also the reason why only at matrix size 8 and input noise precision $10^{-10}$, error deviation is no longer $1$ in Figure 24.
@@ 7.5 Matrix Inversion
Figure 30: Linear correlation between the precision of a matrix determinant (y-axis) to its condition number (x-axis). The legend shows the size of the matrix, as well as the type of the matrix as Random for randomly generated matrix, and Hilbert as the Hilbert matrix.
@@ 8.1 Moving-Window Linear Regression Algorithm
In many applications, data set $Y_{j}$} denotes an input data stream where $j$ represents the time index or sequence index. $Y_{j}$} is thus referred to as a time-series input, with $j$ corresponding to $X_{j}$. A moving window algorithm [13] is applied within a small window centered on each $j$. For each calculation window, $X_{j}=-H,-H+1\dots H-1,H$} where $H$ is an integer constant specifying the half width of the window. This choice ensures $\sum_{j}X_{j}=0$, which simplifies Formula (8.1) and (8.2) into Formula (8.3) and (8.4), respectively [1].
@@
According to Figure LABEL:fig:_Moving_Linear_Fit_Coeff, in which $H$ takes an example value of 2, the values of $(\alpha_{j},\beta_{j})$ can be derived from the previous values $(\alpha_{j-1},\beta_{j-1})$, allowing Formula (8.3) and (8.4) to be reformulated into the progressive moving-window calculation given by Formula (8.5) and (8.6), respectively [1].
@@ 8.2 Variance Adjustment
$\delta^{2}\alpha_{j}$ $=\sum_{X=-H+1}^{H}(\delta Y_{j-H+X})^{2}=\delta^{2}\alpha_{j-1}-(\delta Y_{j-2H})^{2}+(\delta Y_{j})^{2};$ (8.7)
$\delta^{2}\beta_{j}$ $=\sum_{X=-H}^{H}X^{2}(\delta Y_{j-H+X})^{2};$ (8.8)
When the time series contains uncertainty, directly applying Formula (8.5) and (8.6) results in a loss of precision since both formulas reuse each input multiple times, thereby accumulating the variance of that input with every reuse. To prevent this, $\alpha_{j}$ and $\beta_{j}$ should still be calculated progressively using Formula (8.6) and (8.5), respectively, while the variances should instead be computed using Formula (8.7) and (8.8), respectively. Formula (8.8) is not progressive because the progressive form of $\delta^{2}\beta_{j}$ is more expensive in computation than Formula (8.8).
@@
The specified input uncertainty is fixed at $0.2$. Normal noise with a deviation of $0.2$ is added to the slopes, except for the segment $j=10\dots 19$ where Normal noise with a deviation of $2$ is introduced, representing actual uncertainty $10$ times larger than the specified uncertainty.
Figure 31 also presents the results of the moving window fitting of $\alpha+\beta\;Y$ versus the time index $j$. The fitted values of $\alpha$ and $\beta$ follow the expected behave, exhibiting a characteristic delay of $H$ in $j$. When (8.3) and (8.4) are applied to compute the uncertainties of $\alpha$ and $\beta$, both uncertainties increase exponentially with the time index $j$. In contrast, when Formula (8.3) and (8.4) are used exclusively for value calculation, while Formula (8.7) and (8.8) are applied for variance computation, the resulting uncertainties of $\alpha$ and $\beta$ are $\frac{\delta Y}{\sqrt{2H+1}}$, and $\frac{\delta Y}{\sqrt{\frac{H(H+1)(2H+1)}{3}}}$. Both are less than the input uncertainty $\delta Y$, due to the averaging effect of the moving window.
@@ 8.3 Unspecified Input Error
To determine the error deviations of $\alpha$ and $\beta$, the fitting procedure is applied to multiple time-series data sets, each generated with independent noise realizations. Figure 32 illustrates the resulting error deviation as a function of the time index $j$, which remains close to $1$ except within the range $j=10\dots 19$ where the actual noise is ten times greater than the specified value. This observation suggests that an error deviation exceeding 1 may indicate the presence of unspecified additional input errors beyond rounding errors, such as numerical errors in mathematical library functions.
@@ 9.1 DFT (Discrete Fourier Transformation)
For each signal sequence $h[k]$, where $k=0,1\dots N-1$, and $N$ is a natural number, the discrete Fourier transform (DFT) $H[n]$, for $n=0,1\dots N-1$, along with its inverse transformation, is defined by Formula (9.1) and (9.2), respectively [13]. In these expressions, $k$ denotes the time index while $n$ represents the frequency index. The frequency index and time index are not necessarily associated with physical time unit or frequency unit, respectively; rather, the naming convention provides a convenient way to distinguish between the two opposite domains in DFT: the waveform domain $h[k]$ and the frequency domain $H[n]$.
@@
9.2 Modeling Errors of DFT
Figure 33: The DFT spectrum $H[n]$ of signal $h[k]=\sin(f\frac{2\pi}{128}k),k\in[0,127]$, as intensity (y-axis) and phase (embedded y-axis) versus frequency index $n\in[0,18]$ (x-axis and embedded x-axis) for different signal frequency $f$ (legend). This result agrees with both theoretical formula [1] and numerical computation from any mathematical libraries such as SciPy.
Although mathematically self-consistent, DFT implies a periodic boundary condition in the time domain [1]. Consequently, it is only an approximation for the mathematically defined continuous Fourier transform (FT) [1]. For example, the FT spectrum of a sine function is a Delta function at the signal frequency $f$ with a phase $\pi/2$ [13]. Figure 33 shows the DFT spectra of the sine function $h[k]=\sin(f\frac{2\pi}{128}k),k\in[0,127]$, where $f$ is its signal frequency. If DFT is regarded as the digital implementation of FT, the spectra exhibit no modeling error only when the input signal frequency $f$ is an integer, and display varying degrees of modeling errors otherwise. Because of these modeling errors, the use of DFT as the digital implementation of FT is questionable, even though such usage is ubiquitous, and fundamental to many areas of applied mathematics [13].
To avoid the modeling errors inherent in DFT, only Formula (9.1) and (9.2) are used in this paper.
@@ 9.3 FFT (Fast Fourier Transformation)
  • For each output, each input is used only once, therefore no dependency problem arises when decomposing FFT into arithmetic operations as Formula (2.14), (2.15), (2.16), and (2.17).
  • When $L$ is large, the substantial volume of input and output data enables high-quality statistical analysis.
  • The computational complexity is proportional to $L$, since increasing $L$ by 1 adds an additional step involving a sum of multiplications.
  • Each step in the forward transformation doubles the variance, so the uncertainty mean increases with the FFT order $L$ as $\sqrt{2}^{L}$. Because the reverse transformation divides the result by $2^{L}$, its uncertainty mean decreases with $L$ as $\sqrt{1/2}^{L}$. Consequently, the uncertainty mean for the roundtrip transformation is therefore $\sqrt{2}^{L}\times\sqrt{1/2}^{L}=1$.
  • The forward and reverse transformations are identical except for a sign difference, meaning that they are essentially the same algorithm, and any observed difference arises solely from the input data.
@@
  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum in which most values are zeros, causing its uncertainties to grow more rapidly.
  • In contrast, the reverse transformation spreads the precise frequency-domain spectrum (where most values are zeros) back into a time-domain sine or cosine signal, causing its uncertainties to grow more slowly.
@@ 9.4 Testing Signals
  • The results obtained from Sin and Cos signals are statistically indistinguishable.
  • Similarly, the results from Sin and Cos signals at different frequencies does not show significant differences statistically except in the situation of numerical error resonance.
Therefore, the results for Sin and Cos signals across all frequencies are pooled together for statistical analysis, under the unified category Sin/Cos signals.
@@ 9.5 Trigonometric Library Errors
Formula (9.1) and (9.2) restrict the use of $\sin(x)$ and $\cos(x)$ to $x=2\pi j/2^{L}$, where $L$ is the FFT order. To minimize numerical errors in computing $\sin(x)$, the following indexed sine can be used in place of standard library sine functions:
@@
The constructed indexed $\sin(x)$ is referred to as the Quart indexed sine function. In contrast, the direct use of the standard library $\sin(x)$ is referred to as the Library sine function.
Because the the Quart sine function strictly preserves the symmetry and periodicity of sine function, it provides numerical accuracy compared to Library function.
  • Figure 34 shows that the value difference between the Quart and Library $\sin(x)$ and the Quart $\sin(x)$ increases approximately linearly with $|x|$.
  • Figure 35 shows the value difference between the Quart and Library $\cos(x)/\sin(x)$ also increases roughly linearly with $|x|$, but are $10^{2}$ times larger than those observed for $\sin(x)$. Therefore, the linear spectrum in Formula (9.3) may contain significant numerical errors when computed using library sine functions.
For both sine functions, the uncertainty of each $\sin(x)$ is assumed to equal its ULP, which is displayed in Figure 34. Because the Quart sine function has minimal Taylor expansion error due to its minimal range of $x$ in calls to the library $\sin(x)$, its true numerical errors are likely smaller than ULP, therefore its uncertainty is sightly overestimated. In contrast, Figure 34 indicates that the Library sine function exhibits underestimated uncertainties.
@@ 9.6 Using Quart Sine for Sin/Cos Signals
Figure 36: FFT spectrum of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using the Quart sine function after the forward transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
Figure 37: FFT waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using the the Quart sine function after the reverse transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the time index, while the y-axis represents both the uncertainty and the value error.
Figure 38: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal versus FFT order (x-axis) and transformation types (legend) using the Quart sine function.
Figure 39: Histograms of normalized errors of Sin/Cos signals for forward, reverse and roundtrip transformations (legend) using the Quart sine function. The FFT order is $18$.
Using the Quart sine function, for a sine wave with a frequency of $3$, Figure 36 displays the spectrum for the forward transformation, while Figure 37 presents the waveform for the reverse transformation.
  • In the forward transformation, the value errors are slightly smaller than the corresponding result uncertainties, with an error deviation of $0.31$. In the reverse transformation, the error deviation is $0.43$. Both error deviations are less than $1$, confirming a slight overestimation of the sine uncertainty by ULP.
  • The uncertainty mean of the forward transformation is $7$ times larger than that of the reverse transformation. Even with this difference, the uncertainty tracks the corresponding value error effectively.
@@
    As the FFT order increases, the uncertainty mean of the forward transformation grows exponentially faster than that of the reverse transformation. At FFT order $18$, the uncertainty mean of the forward transformation is $10^{3}$ times greater than that of the reverse transformation.
@@
    At FFT order $18$, Figure 39 shows that the normalized error distribution for the reverse transformation is wider than that of the forward transformation, although the two distributions do not differ by an order-of-magnitude. Therefore, the value error increases at least $10^{2}$ times faster in the forward transformation than in the reverse transformation.
@@
    In forward transformation, error deviations reach their stable values rapidly once the FFT order $L\geq 4$. In contrast, in the reverse transformation, error deviations stabilize more slowly with increasing FFT order, approaching stability only when $L\geq 15$. This slow convergence of normalized errors in the reverse transformation is attributed to its input data which consists entirely of precise zeros except at the frequency indexes.
Despite these differences, error deviations are between $0.2$ and $0.5$, suggesting that in variance arithmetic, uncertainties track value errors effectively in all cases, achieving proper coverage.